# Socialized Learning: A Survey of the Paradigm Shift for Edge Intelligence in Networked Systems

### v1

@@ VI-F2 Bidirectional Competition - Struggle
In society, struggles involve complex interactions as people pursue objectives while reacting to others’ actions (as shown in Fig. 36). In inference, layers face bidirectional struggles over shared resources or conflicting goals. Similar to how social struggles inspire innovation, these interactions might lead to optimal solutions. Unmanaged conflicts, however, run the risk of inefficiencies.
  • Resource Allocation: Within EI, resources are finite, and the allocation dynamics mirror social competition for scarce assets. Each layer competes for these resources, driving a push towards greater operational efficiency. However, this competitive environment can also result in increased latency and energy consumption. Collaborative negotiation protocols, such as those proposed by Lee et al. [229], aim to enhance transparent communication and improve efficiency. Concurrently, energy consumption concerns are tackled by methods like those suggested by KN et al. [230], which seek to balance efficiency with assured service quality.
  • Task Offloading: The social dynamics of responsibility-sharing and delegation are reflected in the act of task offloading. Layers wrestle with decisions regarding task delegation, often influenced by their intrinsic capabilities and external pressures. The challenge lies in balancing between local processing and offloading, each with its pros and cons. Nguyen et al. [231] offer an illustrative approach by promoting UAV collaboration, to achieve optimal offloading. Similarly, strategies presented by Wang et al. [232] promote a multi-tiered approach, echoing the social hierarchies and their interactions.
  • Computing Efficiency: Efficiency in computations, much like productivity in society, can be threatened by internal frictions or conflicts. Struggle interactions can give rise to redundant computations or misaligned task allocation. To address these inefficiencies, systems can emulate social cooperation frameworks. Algorithms, as proposed by Xavier et al. [233], mimic social collaboration to ensure optimal resource utilization. Likewise, Amer et al. [234] introduce a hierarchical approach to task scheduling, mirroring social hierarchies, with each level having its own set of responsibilities and interactions.
@@
where $B$ is the level of bidirectional competition, $R(B)$, $T(B)$, and $Comp(B)$ are the efficiency metrics for resource allocation, task offloading, and computation under bidirectional competition, respectively. $\eta_{1}$, $\eta_{2}$, and $\eta_{3}$ are the weights for the respective metrics.
@@
The objective is to fairly and efficiently allocate communication bandwidth and time slots among competing devices, considering their inference task requirements and priorities. Additionally, hierarchical communication resource management frameworks can be employed to coordinate resource allocation across different layers, reducing conflicts and improving overall inference performance. By optimizing communication resource allocation in Struggle interactions, socialized inference systems can achieve more stable and efficient inference.
The bidirectional competition in EI reflects the complex interdependencies familiar in social structures, emphasizing the continuous quest for resources, task offloading efficiency, and computational prowess. Efficiency and complexity are stimulated by the reciprocal rivalry that exists within EI. EI’s competition, which is marked by rivalries at many levels, encourages creativity but makes energy management and communication more challenging. Table XIII describes these features in-depth and emphasizes how they affect the scalability and efficiency of the system.
Figure 37: The Taxonomy Graph of the SL Function.
@@ VI-G Lessons Learned
To ensure smooth interactions within EI, protocols and standardization are necessary, as they align with social norms and regulations. Similar to social issues, protecting sensitive information requires a high priority for security and privacy as EI becomes more interconnected. The efficiency of EI depends on advancing algorithms and technology to address these issues. This technological evolution in EI parallels social advancements, suggesting a future where EI’s complexity is adeptly aligned with the multifaceted nature of society.
@@ VII-A1 Learning
The IEEE’s FL standard IEEE 3652.1-2020 [236], developed by industry leaders, sets a foundational architecture for SL in EI. The importance of this standard lies in its ability to provide a common framework for secure and scalable federated learning, which is a key enabler for SL in EI. By defining guidelines for data governance, model management, and system architecture, IEEE 3652.1-2020 facilitates the deployment of federated learning in various EI applications, such as smart healthcare and industrial IoT. It addresses critical aspects like data, user, and system requirements, enabling secure and scalable SL-driven applications. Further, general standards for privacy and security in IoT and smart cities, like ISO/IEC TS 27570 [237] and ISO/IEC DIS 27400 [238], are important. These standards, along with consumer IoT cybersecurity baseline European Telecommunications Standards Institute (ETSI) EN 303 645 [239], reinforce the framework for trustworthy SL-based EI solutions.
@@ VII-B1 Software
The advancement of EI has pushed the development of a variety of software platforms to simulate and operationalize EI algorithms. Frameworks like TensorFlow Federated and PySyft provide the foundational groundwork for EI simulations. Platforms such as FedML [244], FATE [245], KubeEdge, and HarmonyOS [246] extend these capabilities, offering versatile libraries and secure computing architectures for a broad range of EI applications, from retail to healthcare. For instance, the FedML platform offers a comprehensive framework for federated learning, supporting a wide range of algorithms, models, and datasets. It enables researchers and practitioners to easily develop and deploy federated learning solutions in EI environments, accelerating the adoption of SL in real-world applications. The platform’s modular architecture and extensive documentation make it accessible to users with varying levels of expertise, fostering collaboration and innovation in the field of SL and EI. These platforms play a crucial role in bridging the research-production gap and enhancing real-time EI applications, setting the stage for the integration of SL, which further refines EI through continual user interaction and social inputs.
@@ VII-B2 Solver
In the expanding landscape of EI and wireless networks, resource allocation optimization tools become increasingly critical. Tools like CVX [247], MOSEK [248], and Gurobi [249] are pivotal for large-scale convex optimization and handling mixed-integer conic programs. Techniques like the matrix stuffing technique [250] accelerate the modeling process, enhancing the efficacy of symbolic conic transformation solvers. The integration of these solvers with SL paradigms on EI platforms is essential. An example is the “Open-L2O” framework [251], which blends learning to optimize algorithms, showcasing how SL can be instrumental in adapting EI based on user interactions and feedback.
@@ VIII-A Future Applications of SL in EI
In Section I-C, we provided a concise overview of the current application scenarios of SL in EI, highlighting the common characteristics and potential benefits across various domains. In this section, we expand upon that foundation, offering an in-depth analysis of the future applications of SL in EI. We explore emerging domains and technologies where the integration of SL and EI holds significant promise, and we present a comparative study of these future application scenarios, discussing their unique requirements, challenges, and opportunities. As illustrated in Fig. 38, the integration of SL into EI has an impact on multiple future areas. This section examines various uses and offers guidance on how to put them into practice.
Digital Twins. Digital twins are digital representations of physical entities that simulate real-world scenarios to aid in optimization. Cloud-edge collaborations, e.g., DITEN, enhance this by merging real and virtual systems [255]. Incorporating SL can imbue digital twins with adaptability, allowing them to evolve based on their scenarios. Lu et al.’s DTWN offloads data tasks with digital twins’ help [256]. Sun et al.’s approach uses twins for task offloading; with SL, these twins could offer more adaptive insights [257]. Similarly, a digital twin-enabled space-air network introduces potential for SL in network depictions [258]. In Bellavista et al.’s work on resource allocation, SL guides digital twins in anticipating social needs, enhancing allocation efficiency [259].
Web 3.0. EI and Web 3.0 together drive decentralized applications (DApps) creation on the IoT edge [260]. A key feature of this fusion is real-time interaction through smart contracts and autonomous agents at the network’s edge. Applied to Web 3.0-enhanced EI, SL can amplify the benefits. For instance, integrating SL into DApps development enables systems to adapt dynamically to learned behaviors and interactions, optimizing the user experience. In environments prioritizing data privacy and security, Web 3.0’s decentralization, coupled with SL’s dynamic learning, ensures data security and optimal use in EI [261]. Industrial automation can leverage SL to optimize real-time machine-to-machine interactions, maintaining Web 3.0’s robustness and reliability. Similarly, personalized healthcare can rapidly evolve. With SL-integrated EI in Web 3.0, health applications can learn from individual and global health data, providing tailored, efficient, and timely solutions.
Computing Power Network (CPN). The growth of AI emphasizes the need for an efficient CPN. EI’s integration into CPN promises ultra-low latency and high bandwidth, especially when deployed near the base station [262]. This positioning reduces end-user latency by bridging the gap between content sources and users. However, the union of SL and CPN within EI is what really maximizes its potential. Through SL, the system can dynamically modify computational approaches, learning from node interactions. As industries within the Telecom Network seek customization [263], EI with SL’s algorithms can champion collaborative innovations. Moreover, the ETSI-defined EI platform’s open APIs can harmoniously work with SL, making real-time adjustments from diverse data points. As CPN evolves towards the 6G micro-edge era, SL ensures smooth integrations between EI sites.
Blockchain. The integration of Blockchain and EI has synergized to optimize networking, computing, storage, and security in IoT [264, 265]. Introducing SL to these systems offers tangible benefits. Liang et al. leveraged blockchain with EI for obstacle detection in train autonomous circumambulate system [266]. Incorporating SL here can harness collective insights from blockchain interactions to personalize and enrich the learning process. Sharma et al. emphasized IoT security enhancement through blockchain [267]. SL promotes collaborative security by predicting threats based on historical node interactions. Similarly, Desai et al.’s objective of enhanced device responsiveness in secured systems [268] can benefit from SL, which uses past blockchain transactions to anticipate and mitigate bottlenecks.
AIGC. AIGC uses AI to efficiently produce content, combining expertise and computational power. Integrating SL into AIGC, further enhances its ability to adaptively learn from user interactions and collective insights, making content more relevant and personalized. AIGC, surpassing traditional professionally and user-generated content approaches [269], accelerates content generation, benefiting domains such as text [270], images [271], and videos [272]. A notable example is ChatGPT, which utilizes vast training datasets and operates over 30,000 Nvidia A100 GPUs. The potential of AIGC is magnified when integrated with EI and SL, which address the latency concerns of traditional cloud services. Xu et al. [273] proposed a mobile AIGC framework involving cloud, edge, and D2D mobile layers. In this multi-layered system, the cloud handles the pre-training using generative models like GANs and VAEs, establishing a foundational model. These models are sent to network edges for fine-tuning, aligning with local user needs and enhancing privacy as user preferences remain on-premises. With inferences on proximate edges, latency reduces significantly.
Quantum Information. With the help of qubits, quantum information uses the laws of quantum physics to increase processing capacity, which may be used to raise EI in a variety of ways. For instance, quantum computing can refine AI algorithms in EI by processing vast amounts of data, thus improving predictive accuracy. Quantum computing enhances AI algorithms within EI, thereby improving predictive accuracy and optimizing resources, a benefit especially evident in applications like path planning [274]. This is particularly relevant in evolving quantum AI and in path optimizations. To further fortify EI, quantum communication safeguards edge interactions, establishing unparalleled data security [275].
Metaverse. The Metaverse merges the virtual and real worlds, utilizing technologies like 5G, AI, and VR. It offers users experiences rooted in self-defined identities, socialization, entertainment, and learning [276]. As the complexities of the metaverse grow, there is a trend of moving from cloud to EC to achieve better performance and lower latency. Younis et al. pioneered this by integrating edge-cloud architecture into the metaverse, optimizing bandwidth and ensuring data security [277]. Integrating SL can refine the EI-driven Metaverse. Using SL, the metaverse evolves based on user interactions, fostering an adaptable, virtual environment.
AR and VR. AR and VR revolutionize user interaction with the digital domain. AR blends the real and virtual, while VR immerses users in the virtual. Their need for low-latency interactions is addressed by edge cloud technology, which significantly bolsters their performance [278]. Mobile AR (MAR) emphasizes on-the-go access [279]. Given the intense data requests and latency sensitivity of MAR, hybrid architectures merging cloud, edge, and device components prove effective. In this setup, edges focus on real-time MAR engagement, with clouds handling complex, low-latency tasks [280]. Introducing SL to EI in AR/VR can foster real-time adaptation, refining content based on interactions and feedback.
Comparative Analysis of SL in EI Application Scenarios. While the integration of SL in EI offers significant benefits across various application domains, it is essential to understand the similarities and differences in how SL manifests in these scenarios. We provide a comparative analysis of SL in EI across the key application areas, for example, the similarities in each area are shown below.
  • Collaborative Learning: Across all application scenarios, SL enables collaborative learning among edge devices, allowing them to share knowledge, experiences, and insights to improve overall system performance.
  • Adaptability: SL endows EI systems with the ability to adapt to dynamic environments and user needs by leveraging social principles and collective intelligence.
  • Privacy and Security: The decentralized nature of SL in EI helps protect sensitive data by processing it locally on edge devices, reducing the risk of privacy breaches and security threats.
Accordingly, there are many differences between the different areas of application, as follows.
  • Data Characteristics: The type, volume, and velocity of data vary significantly across application domains. For instance, autonomous driving generates high-volume, real-time sensor data, while IoT systems may deal with more diverse and intermittent data streams.
  • Real-time Requirements: The latency and real-time processing requirements differ across applications. Autonomous driving and AR/VR demand ultra-low latency, while smart city and digital twin applications may tolerate slightly higher latencies.
  • Device Heterogeneity: The diversity and capabilities of edge devices vary across application scenarios. UAVs and autonomous vehicles have more advanced processing capabilities compared to simple IoT sensors or wearable devices in healthcare.
  • SL Objectives: The specific objectives of SL may vary depending on the application. In recommendation systems, SL aims to improve personalization and user satisfaction, while in smart transportation, the focus is on optimizing traffic flow and safety.
Despite these differences, the core principles of SL, such as collaborative learning, adaptability, and privacy preservation, remain consistent across all application domains. The specific implementation of SL algorithms and architectures may need to be tailored to the unique requirements and constraints of each scenario. For example, SL algorithms for autonomous driving may prioritize low-latency communication and real-time decision-making, while those for smart healthcare may emphasize privacy-preserving data analytics and personalized treatment recommendations.

### v2

@@ VI-F2 Bidirectional Competition - Struggle
In society, struggles involve complex interactions as people pursue objectives while reacting to others’ actions (as shown in Fig. 36). In inference, layers face bidirectional struggles over shared resources or conflicting goals. Similar to how social struggles inspire innovation, these interactions might lead to optimal solutions. However, unmanaged conflicts run the risk of inefficiencies.
  • Resource Allocation: Within EI, resources are finite, and the allocation dynamics mirror the social struggle for scarce resources. Each layer competes for these resources, driving a push towards greater operational efficiency. However, this struggle can also result in increased latency and energy consumption. Collaborative communication protocols, such as those proposed by Lee et al. [232], aim to enhance transparent communication and improve efficiency. Concurrently, energy consumption concerns are tackled by methods like those suggested by KN et al. [233], which seek to balance efficiency with assured QoS.
  • Task Offloading: The social dynamics of responsibility-sharing and allocation is reflected in the act of task offloading. Layers struggle with decisions regarding task offloading, often influenced by their intrinsic capabilities and external pressures. The challenge lies in balancing between local processing and offloading, each with its pros and cons. Nguyen et al. [234] offer an illustrative approach by promoting UAV collaboration, to achieve optimal offloading. Similarly, strategies presented by Wang et al. [235] promote a multi-tiered approach, echoing the social hierarchies and their interactions.
  • Computing Efficiency: Efficiency in computations, much like productivity in society, can be threatened by internal frictions or conflicts. Struggle interactions can give rise to redundant computations or misaligned task allocation. To address these issues, systems can emulate social cooperation frameworks. Algorithms, as proposed by Xavier et al. [236], mimic social cooperation to ensure optimal resource utilization. Amer et al. [237] introduce a hierarchical approach to task scheduling, mirroring social hierarchies, with each level having its own set of responsibilities and interactions.
@@
where $B$ is the level of bidirectional struggle, $R(B)$, $T(B)$, and $Comp(B)$ are the efficiency metrics for resource allocation, task offloading, and computation under bidirectional struggle, respectively. $\eta_{1}$, $\eta_{2}$, and $\eta_{3}$ are the weights for the respective metrics.
@@
The objective is to fairly and efficiently allocate communication bandwidth and time slots among struggle devices, considering their inference task requirements and priorities. Additionally, hierarchical communication resource management frameworks can be employed to coordinate resource allocation across different layers, reducing conflicts and improving overall inference performance. By optimizing communication resource allocation in struggle interactions, socialized inference systems can achieve more stable and efficient inference.
The bidirectional struggle in EI reflects the complex interdependencies familiar in society, emphasizing the continuous pursuit of resources, task offloading efficiency, and computing power. Efficiency and complexity are stimulated by the struggle that exists within EI. EI’s struggle, which is marked by rivalries at many levels, encourages creativity but makes energy management and communication more challenging. Table XIII describes conflict interactions in-depth and emphasizes how they affect the scalability and efficiency of the system.
Figure 37: The taxonomy graph of the unified SL function. It maps the diverse methods of socialized training and inference in EI to the variables $(\mathbf{A},\mathbf{R}_{w},\mathbf{R}_{a},\mathbf{I},\mathbf{L})$ defined in Eq. (1). This enables researchers to design customized and optimized SL schemes by selecting appropriate values for these variables, tailored to specific EI application requirements.
@@ VI-G Lessons Learned
To ensure seamless interactions within EI, protocols and standardization are necessary, as they align with social norms and regulations. Similar to social issues, protecting sensitive information requires a high priority for security and privacy as EI becomes more interconnected. The efficiency of EI depends on advancing algorithms and technology to address these issues. This technological evolution in EI parallels social advancements, suggesting a future where EI’s complexity is adeptly aligned with the multifaceted nature of society.
@@ VII-A1 Learning
The IEEE’s FL standard IEEE 3652.1-2020 [239], developed by industry leaders, sets a foundational architecture for SL in EI. The importance of this standard lies in its ability to provide a common framework for secure and scalable FL, which is a key enabler for SL in EI. By defining guidelines for data governance, model management and system architecture, IEEE 3652.1-2020 facilitates deployment of FL in various EI applications, such as smart healthcare and industrial IoT. It addresses critical aspects like data, user, and system requirements, enabling secure and scalable SL-driven applications. Further, general standards for privacy and security in IoT and smart cities, like ISO/IEC TS 27570 [240] and ISO/IEC DIS 27400 [241], are important. These standards, along with consumer IoT cybersecurity baseline European Telecommunications Standards Institute (ETSI) EN 303 645 [242], reinforce framework for trustworthy SL solutions in EI.
@@ VII-B1 Software
The advancement of EI has pushed the development of a variety of software platforms to simulate and operationalize EI algorithms. Frameworks like TensorFlow Federated and PySyft provide the foundational groundwork for EI simulations. Platforms such as FedML [246], FATE [247], KubeEdge, and HarmonyOS [248] extend these capabilities, offering versatile libraries and secure computing architectures for a broad range of EI applications, from retail to healthcare. For instance, the FedML platform offers a comprehensive framework for FL, supporting a wide range of algorithms, models, and datasets. It enables researchers and practitioners to easily develop and deploy FL solutions in EI environments, accelerating the adoption of SL in real-world applications. The platform’s modular architecture and extensive documentation make it accessible to users with varying levels of expertise, fostering collaboration and innovation in the field of SL and EI. These platforms play a crucial role in bridging the research-production gap and enhancing real-time EI applications, setting the stage for the integration of SL, which further refines EI through continual user interaction and social inputs.
Figure 38: Future applications of SL in EI.
@@ VII-B2 Solver
In the expanding landscape of EI and wireless networks, resource allocation optimization tools become increasingly critical. Tools like CVX [249], MOSEK [250], and Gurobi [251] are pivotal for large-scale convex optimization and handling mixed-integer conic programs. Techniques like the matrix stuffing technique [252] accelerate the modeling process, enhancing the efficiency of symbolic conic transformation solvers. The integration of these solvers with SL paradigms on EI platforms is essential. An example is the “Open-L2O” framework [253], which blends learning to optimize algorithms, showcasing how SL can be instrumental in adapting EI based on user interactions and feedback.
@@ VIII-A Future Applications of SL in EI
In Section I-C, we provided a concise overview of the current application scenarios of SL in EI, highlighting the common characteristics and potential benefits across various domains. This section provides an in-depth analysis of SL in EI’s future applications, exploring emerging domains and technologies. We comparatively study these application scenarios, discussing their unique requirements, challenges, and opportunities. As illustrated in Fig. 38, the integration of SL into EI has an impact on multiple future application scenarios. This section examines various uses and offers guidance on how to put them into practice.
Digital Twins. Digital twins, as digital representations of physical entities, are evolving to incorporate SL for enhanced adaptability and optimization. Implementing complex digital twin models efficiently on resource-constrained edge devices is a key challenge. Future research should focus on developing lightweight, yet high-precision physical models and distributed training techniques. For instance, Lu et al.’s DTWN [257] could be extended to incorporate FL, enabling collaborative model updates across multiple edge nodes while preserving data privacy. Sun et al.’s task offloading approach [258] could be enhanced with RL-based decision-making to dynamically optimize offloading strategies based on network conditions and task characteristics. Future work in space-air networks [259] could explore hierarchical digital twins leveraging SL for multi-level optimization from individual device management to global network orchestration. Bellavista et al.’s resource allocation research [260] could be expanded to include predictive analytics, utilizing SL to anticipate and proactively adjust to resource demands.
Web 3.0. The integration of EI and Web 3.0 promotes the creation of decentralized applications (DApps) on IoT edges [261], which poses unique challenges for SL. Developing efficient consensus mechanisms in decentralized environments is a critical issue. Future research should investigate SL-enhanced consensus algorithms, like reputation-based systems where node credibility hinges on SL performance. Integrating SL in DApp development could enable adaptive smart contracts that evolve based on user interactions and network conditions. This necessitates novel contract languages that handle probabilistic and learning-based clauses. Privacy-preserving SL techniques require optimization on resource-constrained edge devices to ensure data security in EI environments [262]. Research in industrial automation should focus on real-time SL algorithms for machine-to-machine interactions, employing online learning techniques to adapt to changing production conditions while upholding the robustness of Web 3.0.
Computing Power Network (CPN). Integrating EI into CPN presents optimization challenges for achieving ultra-low latency and high bandwidth, especially near base stations [263]. Future studies should develop dynamic resource allocation algorithms that adjust in real time to SL demands. This approach requires predictive resource allocation techniques that utilize learned SL task patterns. Designing distributed learning algorithms that efficiently utilize heterogeneous computing resources across the CPN is a key challenge. As industries demand customization [264], research must explore adaptive SL algorithms that dynamically adjust to varying computing resources and network conditions. The shift towards 6G micro-edge architectures demands novel SL frameworks that operate seamlessly across diverse EI sites, leveraging FL and others to balance data locality with global model optimization.
Blockchain. Although the integration of Blockchain and EI optimizes IoT systems [49, 265], it presents challenges in efficiency and scalability. Future studies should develop SL-enhanced consensus mechanisms tailored for resource-constrained edge devices. Liang et al.’s work on obstacle detection in train systems [266] could adopt FL to enable collaborative model updates while preserving data privacy across distributed nodes. Designing SL algorithms that leverage blockchain’s immutability for secure, decentralized learning presents a key challenge. Building on Sharma et al.’s IoT security framework [267], future research should explore using distributed SL in blockchain-based anomaly detection to predict threats from historical node interactions. To enhance scalability in Desai et al.’s system [268], research should focus on combining sharding techniques with SL to boost throughput in large-scale IoT networks.
AIGC. Integrating SL into AIGC enhances its ability to adaptively learn from user interactions and collective insights, personalizing content, but also introduces unique challenges beyond traditional approaches [269]. Future research should focus on SL-enhanced generative models that adapt to collective user behaviors at the edge. Building on Xu et al.’s multi-layered framework [270], developing SL algorithms that facilitate collaborative learning of GANs and VAEs across distributed edge nodes is necessary. Designing privacy-preserving SL techniques for personalized content generation, which potentially leverage FL with differential privacy, is a critical challenge. Research should focus on exploring SL-driven adaptive model partitioning that dynamically allocates tasks between different layers, based on social patterns and device capabilities. Furthermore, investigating SL-based content optimization techniques at the edge for various modalities, such as text [271], images [272], and videos [273], while balancing quality, latency and resource constraints, is crucial.
Quantum Information. Integrating quantum computing principles with SL in EI presents unique challenges for enhancing distributed learning and security. Research should prioritize developing quantum-inspired SL algorithms to enhance collaborative learning efficiency in edge networks. For path planning [274], quantum-inspired optimization techniques could enable edge devices to collaboratively solve complex problems more efficiently. Designing privacy-preserving SL protocols that leverage quantum-resistant cryptography to ensure long-term security against advancing quantum capabilities remains a key challenge. Investigating quantum key distribution (QKD) for securing SL data exchanges in edge networks is crucial [275]. Future work should aim to integrate quantum-safe cryptography within SL frameworks, focusing on scalability and resilience in real-world edge environments.
Metaverse. The Metaverse, which merges virtual and real worlds through 5G, AI, and VR technologies, presents unique challenges for EI and SL integration [276]. As Metaverse complexity grows, the trend towards EC becomes crucial for improving performance and reducing latency. Extending Younis et al.’s edge-cloud architecture [277], future research should focus on SL-enhanced bandwidth optimization and data security in Metaverse environments. Key challenges include developing adaptive SL algorithms for dynamic edge-cloud resource allocation based on real-time user interactions and network conditions. Research into SL techniques that efficiently synchronize heterogeneous data types (visual, auditory, haptic) across distributed edge nodes in the Metaverse is needed. Furthermore, exploring SL-driven content adaptation mechanisms is crucial for developing virtual environments that evolve in response to collective user behaviors.
AR and VR. AR/VR require low-latency interactions, which are facilitated by edge-cloud technologies [278]. Future studies should develop SL-enhanced task allocation algorithms for hybrid cloud-edge-device architectures, focusing on Mobile AR [279]. Designing adaptive SL models that dynamically allocate computational tasks between edge and cloud, responding to real-time user interactions and network conditions, is a major challenge. Given intense data demands and latency sensitivity, investigating SL techniques for predictive content caching and rendering at the edge is crucial [280]. Research into collaborative SL algorithms that facilitate real-time content adaptation across multiple AR/VR devices, leveraging collective user interactions to create personalized experiences, is needed. Furthermore, exploring SL-driven compression techniques for data streams could reduce bandwidth requirements while preserving QoE.
Comparative Analysis of SL in EI Application Scenarios. Integrating SL in EI across various domains highlights both commonalities and unique challenges. While collaborative learning, adaptability, privacy preservation and others are core principles of SL in EI, their implementation varies significantly depending on application-specific requirements. Data characteristics and real-time processing demands vary significantly across domains: digital twins and metaverse applications process complex, multi-modal data streams, whereas Web 3.0 and blockchain systems handle distributed, often encrypted data. This variability necessitates developing adaptive SL frameworks that can efficiently handle heterogeneous data types and processing requirements.
Device heterogeneity presents another critical challenge, as the computational capabilities range from ESs in CPNs to resource-constrained IoT devices in AIGC scenarios. Future research should develop scalable SL algorithms capable of dynamically adjusting to diverse device capabilities, utilizing techniques like model compression, adaptive quantization, and intelligent task offloading. The objectives of SL also differ across applications: AR/VR systems prioritize low-latency content adaptation, whereas quantum-enhanced EI focuses on secure, distributed computation.
Despite these differences, it is crucial to maintain SL’s core principles across all applications. Future research should investigate unified SL frameworks that can adapt to diverse scenarios while ensuring collaborative learning, system adaptability, and privacy preservation. This approach could involve developing modular SL architectures with interchangeable components tailored to various application requirements, or exploring meta learning techniques that rapidly adapt to new domains with minimum fine-tuning. By addressing these challenges, SL in EI can evolve to meet the diverse needs of future applications while maintaining its fundamental advantages in collaborative, adaptive, and secure distributed learning.